# 3DSRBench: A Comprehensive 3D Spatial Reasoning Benchmark

### v3

@@ 4.3 Robustness to Uncommon Camera Viewpoints
Figure 5: Failures cases of GPT-4o [29] (top) and Gemini 2.0 Flash thinking [23] (bottom) on our 3DSRBench. (a) GPT-4o: GPT-4o does not have an explicit 3D representation, e.g., metric depth, and resort to visual cues to compare the distance, which leads to a wrong answer. (b) Gemini 2.0 Flash thinking: In this example Gemini 2.0 Flash thinking successfully breaks down the 3D spatial reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning and predicts a wrong answer. See Sec. E in supplementary materials for more failure cases of the two models.
@@
As demonstrated by the results in Tab. 4, all LMMs exhibit significantly degraded performance when generalizing from common to uncommon viewpoints, e.g., a 13.5% drop in accuracy for GPT-4o [29], a 32.2% drop for Gemini-1.5-Pro [51], and a 19.1% drop for LLaVA-NeXT-8B [40]. We visualize two failure examples of GPT-4o in Fig. 3, showing how it is capable of predicting the correct answer when prompting with an image from a common viewpoint but fails when asked the same question with an image rendered from an uncommon viewpoint of the exact same scene.
@@
These findings show that 3D spatial reasoning abilities of state-of-the-art LMMs are not robust to uncommon camera viewpoints. This largely limits their applicability to various downstream applications in robotics and embodied AI. Cameras mounted on robot arms or embodied AI systems are often positioned in uncommon locations and orientations as used in our study (see Fig. 3). On the one hand impressive advancements achieved by state-of-the-art LMMs in standard spatial reasoning benchmarks [30, 28, 36] may not generalize to downstream tasks; on the other hand, significantly degraded performance in uncommon viewpoints raises serious concerns about AI safety [3].
@@ 4.4 Failure Cases
We present two failure cases of GPT-4o [29] and Gemini 2.0 Flash thinking [23] in Fig. 5. In Fig. 5(a) we see that GPT-4o cannot perform rigorous 3D spatial reasoning and resort to various visual cues for reasoning. This is because GPT-4o lacks explicit 3D representations, e.g., metric depth, that limits its ability to perform complex 3D spatial reasoning. In Fig. 5(b), Gemini 2.0 Flash thinking successfully breaks down the 3D reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning step-by-step. Despite the good thinking, the model fails to follow the planning and predicts a wrong answer.
We argue that for 3D spatial reasoning problems, models must not only have strong visual encoders to parse 3D-aware features, but also build a powerful reasoning model on various 3D information. Although scaling language model size leads to stronger reasoning abilities (see Fig. 4), a lack of explicit 3D representations would fundamentally limit models’ abilities to solve complex 3D spatial reasoning questions that require multi-step 3D computations.
@@ C Qualitative Examples of 3DSRBench
We present two example questions for each of the 12 question types in Fig. 6 (height and location questions), Fig. 7 (orientation questions), and Fig. 8 (multi-object reasoning questions).
@@ D Qualitative Examples of Common and Uncommon Viewpoints
We present some qualitative examples of 3DSRBench-synthetic with multi-view images rendering the same scene and objects but from common (left) and uncommon (right) viewpoints in Fig. 9.
@@ E Failure Cases
We present some failures cases of GPT-4o [29] in Fig. 10 and of Gemini 2.0 Flash thinking [23] in Fig. 11.

### v4

@@ 4.3 Robustness to Uncommon Camera Viewpoints
Figure 5: Failures cases of GPT-4o [28] (top) and Gemini 2.0 Flash thinking [22] (bottom) on our 3DSRBench. (a) GPT-4o: GPT-4o does not have an explicit 3D representation, e.g., metric depth, and resort to visual cues to compare the distance, which leads to a wrong answer. (b) Gemini 2.0 Flash thinking: In this example Gemini 2.0 Flash thinking successfully breaks down the 3D spatial reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning and predicts a wrong answer. See Sec.˜E in supplementary materials for more failure cases of the two models.
@@
As demonstrated by the results in Tab.˜4, all LMMs exhibit significantly degraded performance when generalizing from common to uncommon viewpoints, e.g., a 13.5% drop in accuracy for GPT-4o [28], a 32.2% drop for Gemini-1.5-Pro [53], and a 19.1% drop for LLaVA-NeXT-8B [39]. We visualize two failure examples of GPT-4o in Fig.˜3, showing how it is capable of predicting the correct answer when prompting with an image from a common viewpoint but fails when asked the same question with an image rendered from an uncommon viewpoint of the exact same scene.
@@
These findings show that 3D spatial reasoning abilities of state-of-the-art LMMs are not robust to uncommon camera viewpoints. This largely limits their applicability to various downstream applications in robotics and embodied AI. Cameras mounted on robot arms or embodied AI systems are often positioned in uncommon locations and orientations as used in our study (see Fig.˜3). On the one hand impressive advancements achieved by state-of-the-art LMMs in standard spatial reasoning benchmarks [29, 27, 35] may not generalize to downstream tasks; on the other hand, significantly degraded performance in uncommon viewpoints raises serious concerns about AI safety [3].
@@ 4.4 Failure Cases
We present two failure cases of GPT-4o [28] and Gemini 2.0 Flash thinking [22] in Fig.˜5. In Fig.˜5(a) we see that GPT-4o cannot perform rigorous 3D spatial reasoning and resort to various visual cues for reasoning. This is because GPT-4o lacks explicit 3D representations, e.g., metric depth, that limits its ability to perform complex 3D spatial reasoning. In Fig.˜5(b), Gemini 2.0 Flash thinking successfully breaks down the 3D reasoning question into small and tractable steps. However, without explicit 3D representations, the model cannot perform reliable 3D spatial reasoning step-by-step. Despite the good thinking, the model fails to follow the planning and predicts a wrong answer.
We argue that for 3D spatial reasoning problems, models must not only have strong visual encoders to parse 3D-aware features, but also build a powerful reasoning model on various 3D information. Although scaling language model size leads to stronger reasoning abilities (see Fig.˜4), a lack of explicit 3D representations would fundamentally limit models’ abilities to solve complex 3D spatial reasoning questions that require multi-step 3D computations.
@@ 5 Conclusions
Acknowledgements
We would like to thank Yiyan Li, Lizhi Ma, and the anonymous reviewers for their helpful comments and suggestions. Wufei Ma and Alan Yuille acknowledges support from ONR with N00014-23-1-2641 and ARL award with W911NF2320008.
@@ C Qualitative Examples of 3DSRBench
We present two example questions for each of the 12 question types in Fig.˜6 (height and location questions), Fig.˜7 (orientation questions), and Fig.˜8 (multi-object reasoning questions).
@@ D Qualitative Examples of Common and Uncommon Viewpoints
We present some qualitative examples of 3DSRBench-synthetic with multi-view images rendering the same scene and objects but from common (left) and uncommon (right) viewpoints in Fig.˜9.
@@ E Failure Cases
We present some failures cases of GPT-4o [28] in Fig.˜10 and of Gemini 2.0 Flash thinking [22] in Fig.˜11.